OpenStudy (anonymous):

How do you do this equation 4b+3=-9

OpenStudy (mayankdevnani):

4b=-9-3 4b=-12 b=-12/4=-3

OpenStudy (mayankdevnani):

ok @tatetate

OpenStudy (anonymous):

Like how is it writtten out please exuse my stupitdy im only n 7th grade .

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

Like do i write it out the wy you just put it?

OpenStudy (mayankdevnani):

thnx

OpenStudy (anonymous):

@mayankdevnani dont just hand out the answers

OpenStudy (mayankdevnani):

ok